OpenStudy (anonymous):

x+2 3 ----- = -- x-2 2

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

(x + 2)2 = 3(x-2) 2x+4 = 3x-6 4+6 = 3x-2x 10 = x x = 10